# ChatEMG: Synthetic Data Generation to Control a Robotic Hand Orthosis for Stroke

### v1

@@ Abstract
Intent inferral on a hand orthosis for stroke patients is challenging due to the difficulty of data collection from impaired subjects. Additionally, EMG signals exhibit significant variations across different conditions, sessions, and subjects, making it hard for classifiers to generalize. Traditional approaches require a large labeled dataset from the new condition, session, or subject to train intent classifiers; however, this data collection process is burdensome and time-consuming. In this paper, we propose ChatEMG, an autoregressive generative model that can generate synthetic EMG signals conditioned on prompts (i.e., a given sequence of EMG signals). ChatEMG enables us to collect only a small dataset from the new condition, session, or subject and expand it with synthetic samples conditioned on prompts from this new context. ChatEMG leverages a vast repository of previous data via generative training while still remaining context-specific via prompting. Our experiments show that these synthetic samples are classifier-agnostic and can improve intent inferral accuracy for different types of classifiers. We demonstrate that our complete approach can be integrated into a single patient session, including the use of the classifier for functional orthosis-assisted tasks. To the best of our knowledge, this is the first time an intent classifier trained partially on synthetic data has been deployed for functional control of an orthosis by a stroke survivor.
@@ I Introduction
Wearable rehabilitation robots continuously interact with human patients and must make constant decisions on when and how to provide motor assistance. Machine learning methods are increasingly being applied to mediate this stream of information. However, many current learning methods, which have revolutionized domains such as vision or language understanding, are predicated on the availability of large training datasets. Compared to these domains, applications for learning on wearable robots are faced with a tremendous scarcity of both raw data and reliable ground truth labels.
One such example is intent inferral, or the process by which a robotic orthosis or prosthesis collects a set of biosignals from the user, and uses them to infer the activity that the user intends to perform, so it can provide the right type of physical assistance at the right moment. In the case of a hand orthosis for stroke survivors developed in our lab [1, 2] (Fig. 1), the device can use forearm electromyographic (EMG) data to predict when the user is trying to open the hand, and provide assistance in order to overcome the muscle spasticity that is a common stroke after-effect. Widely considered to be a key problem in assistive and rehabilitative robotics [3], an effective intent inferral mechanism can be an intuitive way to control a robotic device, avoiding cognitive overload and enabling long-term use.
@@
Like many other applications in the assistive and rehabilitation devices domain, a fundamental challenge in intent inferral for stroke is the difficulty of collecting training data. The variation in EMG signals among conditions, sessions, and subjects makes the challenge even more pronounced. There are significant variations of EMG signals among different stroke survivors due to variations in neuromuscular control impairments. Defining a use session as a single use of the device between donning and doffing, we note that even for the same subject, the muscle tone and spasticity can vary across different sessions. Furthermore, the signals could change within a single session, depending on the use conditions, such as the hand position and whether the motor is engaged and providing active grasping assistance [4, 5]. Due to such variation, intent classifiers trained for a specific condition, session, or subject, do not generalize well, and classical solutions often require tedious data collection on every new condition, session, or subject. This introduces a significant burden on the participants.
In this work, we aim to reduce the burden of data collection from stroke subjects by generating synthetic data with generative models. We propose ChatEMG, an autoregressive generative model that understands the broad behavior of forearm EMG signals from a corpus of offline data across different stroke subjects and then can generate personalized (i.e., condition-, session-, and subject-specific) synthetic samples conditioned on prompts (i.e., a given sequence of EMG signals) sampled from a very limited dataset of a new condition, session, or subject.
@@ II Related Work
There are many previous works that attempt to infer the activity intents of disabled-bodied subjects using EMG signals with machine learning. These predicted intents are often sent to control wearable assistive and rehabilitative devices. A majority of works [4, 6, 7, 8] use supervised learning to control prostheses or orthoses, where the model is trained only on an initial dataset and then used during a longer session. However, it is not easy to collect labeled data on stroke subjects, and a new dataset needs to be collected for every new session after donning and doffing the device. Recent works [5, 9, 10, 11, 12, 13, 14, 15, 16] have been exploring the semi-supervised learning paradigm to control wearable devices with EMG signals. This paradigm leverages unlabeled data to make the intent inferral algorithm more robust to change from the input signals within a single session. It still relies on some heuristics to “label” the unlabeled data, which can be inaccurate from time to time.
@@
The lack of accurate and reliable data is not specific to intent inferral, but to the general machine learning research in the medical community. Similar to ours, there is some previous work that studies generative models for synthetic data generation. A majority of them use Generative Adversarial Network (GAN) [17] and its variants. Shin et al. [18] use GAN to generate synthetic MRI images to detect brain tumors. Hazra et al. [19] propose SynSigGAN with a novel preprocessing stage that generates four types of biomedical signals, including chin EMG signals during sleep for able-bodied subjects. Zanini et al. [20] use style transfer and DCGANs to generate EMG signals for Parkinson’s Disease. Pinto et al. [21] use WGAN-GP to generate EMG signals for six basic hand gestures for healthy subjects.
GAN and its variants are not autoregressive and can only generate a fixed length of EMG signals from random seeds. In comparison, ChatEMG can generate an unlimited sequence of EMG signals conditioned on EMG prompts, which are also of arbitrary length. This is extremely useful in applications where personalization is necessary. After stroke, each subject might have different hand functionality or muscle coactivation patterns, and ChatEMG allows us to bias data generation via such prompting. Most similar to ours, Bird et al. [22] train a GPT-2 model to generate EMG signals for hand open/close classification for healthy subjects. However, we specifically focus here on stroke patients whose gestures are much more challenging to infer due to abnormal muscle activation.
@@ III Overview
The most direct way to achieve this goal is to collect a set of training data (EMG signals) labeled with ground truth intent. This can be done by instructing the patient to attempt one of the three hand movements of interest, while simultaneously recording EMG data and labeling it with the prescribed intent. Once enough labeled data is collected for each of the three possible intent classes, we can train a classifier to distinguish between them. Numerous types of models are available for such a classifier, ranging from linear classifiers and decision trees to neural networks.
However, this traditional approach suffers from two key limitations: (1) The process of collecting labeled training data is burdensome and time-consuming for both the patient and the experimenter. It uses up precious session time and also fatigues the patients, leading to increased muscle tone and spasticity. (2) EMG signals exhibit significant variations between different conditions, sessions, and subjects. This means that training data collected in one condition/session/subject is unlikely to apply to a different one, leading to very poor generalization performance of the classifier. Collecting new training data for every new condition/session/subject makes the difficulty of data collection even more pronounced.
@@ IV ChatEMG Generative Models
The role of a ChatEMG model is to take in a sequence of EMG signals as input and predict the next EMG signal as output, where one signal consists of the 8-channel data from our EMG armband. Such a model can then be used autoregressively to generate synthetic data of arbitrary length, all conditioned on the given prompt. We note that this approach is similar in concept to language models such as ChatGPT [23], capable of generating text in response to a text prompt. However, ChatEMG operates on the “language” of EMG data, hence its chosen moniker.
As noted, we train one ChatEMG model on data corresponding to each user intent. The goal is then for each of these models to generate synthetic data corresponding to the intent on which it has been trained. However, while the data on which each of these models is trained is different, their architecture is the same, as described below.
@@ IV-A Architecture
One key consideration for designing the ChatEMG architecture is its ability to generate diverse output given the same input signals. As a result, we use an architecture where the output of the model is sampled from a probability distribution over a discrete vocabulary. Furthermore, we can generate an arbitrary number of “likely” next signals simply by repeating the sampling process. Each sequence can then be continued in an autoregressive fashion. This means that one generative model can use a single prompt in order to generate an arbitrary number of likely next sequences, each of them of arbitrary length. This is an important feature of our model, given that we will need large amounts of synthetic data in order to train downstream classifiers [^1] As an alternative, we also tried a variational regression setup that predicts a mean and variance for the next signal, and used the reparameterization trick to backpropagate the loss, which also allows sampling an arbitrary number of completions for a single prompt. However, we found empirically that the balance between the KL (Kullback-Leibler) loss and the reconstruction loss is hard to tune, and its generated signals are not as good as the classification version..
However, this architecture requires a discrete vocabulary of finite size, which is challenging given that our raw data consists of an 8-channel EMG signal. For preprocessing, we bin and clip each channel of the EMG signal to be an integer between 0 and 1000. Under this range, if we model the whole 8-channel signal as one “token” in our vocabulary, then our vocabulary size becomes $1000^{8}$, which is too large to predict a probability distribution and sample from. As a result, ChatEMG always predicts the next EMG value for only the first channel, making the vocabulary size 1000, which is feasible to predict a probability distribution over.
@@ IV-B Modelling Inter-channel Relationship
Despite the fact that our model predicts one channel at a time, ChatEMG still considers the interchannel relationship across different channels. Our EMG armband has 8 electrodes surrounding the forearm, and each electrode covers a particular area of the forearm muscle. The interchannel relationship across different electrodes can be useful information to leverage for understanding the user’s intent.
@@
Our model is trained with EMG sequences of length $t=256$, which corresponds to 2.56 seconds as our armband collects at 100Hz. The dimension of the model output is $256\times 1$, and they are the 256 predicted EMG values for the next time step of the first channel. During training, we also augment the input signals by rotating the channels seven times to simulate the rotation of the electrodes. This data augmentation strategy enables ChatEMG to be invariant with channel rotation. During data generation, we sample EMG prompts of length 150 (corresponding to 1.5s) from the very limited dataset of the new condition, session, or subject and use ChatEMG to autoregressively complete the rest of the signal to a length of 256, which is the time-series length that our classification algorithms take.
@@ V Experimental Setup
The ultimate goal of the orthosis is to provide meaningful functional assistance for stroke subjects, enabled by intent inferral. Thus, we evaluate ChatEMG by examining if the generated synthetic samples can improve intent inferral accuracy for stroke participants.
@@ V-A Subjects
We performed experiments with five chronic stroke survivors having hemiparesis and moderate muscle tone: Modified Ashworth Scale (MAS) scores $\leq$ 2 in the upper extremity. Our MAS criteria exclude subjects whose fingers are difficult to move passively — fingers with more severe spasticity cannot be quickly extended with external forces without increasing muscle tone and risking damage to the joints. Our participants can fully close their hands but are unable to completely extend their fingers without assistance. The passive range of motion in the fingers is within functional limits. Testing was approved by the Columbia University Institutional Review Board (IRB-AAAS8104) and was performed in a clinical setting under the supervision of an occupational therapist.
@@ V-B Data Collection Protocol
For each stoke subject, we collect data from two sessions on two different days (at least a week apart), with the following protocol. A session is defined to be a single use of the device between donning and doffing the device. We intentionally keep the two sessions for each subject at least one week apart to better study the variation of the EMG signals across different days. For each session, we collect data under four different conditions: 1) with the arm resting on a table and the orthosis motor off {arm on table, motor off}, 2) with the arm resting on a table and the orthosis motor on, providing active grasp assistance {arm on table, motor on}, 3) with the arm raised above the table and the orthosis motor off {arm off table, motor off}, and 4) with the arm raised above the table and the orthosis motor on {arm off table, motor on}.
@@ V-C Assessment Scenarios
We create different assessment scenarios that simulate different use cases of ChatEMG, by selecting different training recordings (recordings used to train ChatEMG) and intent inferral recordings (recordings used to perform intent inferral evaluation). These scenarios evaluate how well ChatEMG generalizes and adapts to new conditions, sessions, and subjects not seen in its training recordings. The detailed descriptions are as follows.
@@ V-C3 Subject Adaptation
This scenario seeks to address the challenge of onboarding new subjects. We conduct five separate experiments, each one simulating the onboarding of one holdout subject, given that we have seen the other four. In each experiment, we train ChatEMG using all the recordings from the other subjects (including both sessions), and the intent inferral recordings are those of the holdout subject. When adapting to a new subject, it is also implicitly adapting to a new session. However, in our session adaption experiments, we assume it is a different session of a previously seen subject.
@@ V-D Intent Inferral Classifiers
ChatEMG is classifier-agnostic, and the generated synthetic samples can be integrated with the training set of any classifiers. We study three types of classifiers: linear discriminant analysis (LDA), random forests (RF), and Transformer. They are popular in the biomedical literature and cover both classic machine-learning algorithms and high-capacity neural networks. We feed into each classifier a time series of length 256 (2.56s), in the shape of 256 by 8. The EMG signals are flattened into a single vector for LDA and RF. We use a single 4-head attention block followed by a 3-layer multilayer perception (MLP) for the Transformer classifier.
@@ V-E Baselines
The intent inferral evaluation is done on individual intent inferral recordings. For each recording, we assume only a small support set (i.e., the first open-relax-open motion of the recording) is available for training the classifier. The support set simulates the limited new training samples from a new condition, session, or subject, denoted by $\mathcal{D}_{\text{new}}^{\text{synth}}$ in Sec. III. We then test the classifier’s accuracy using the query set (i.e., the second and third open-relax-close motions).

### v3

@@ Abstract
Intent inferral on a hand orthosis for stroke patients is challenging due to the difficulty of data collection. Additionally, EMG signals exhibit significant variations across different conditions, sessions, and subjects, making it hard for classifiers to generalize. Traditional approaches require a large labeled dataset from the new condition, session, or subject to train intent classifiers; however, this data collection process is burdensome and time-consuming. In this paper, we propose ChatEMG, an autoregressive generative model that can generate synthetic EMG signals conditioned on prompts (i.e., a given sequence of EMG signals). ChatEMG enables us to collect only a small dataset from the new condition, session, or subject and expand it with synthetic samples conditioned on prompts from this new context. ChatEMG leverages a vast repository of previous data via generative training while still remaining context-specific via prompting. Our experiments show that these synthetic samples are classifier-agnostic and can improve intent inferral accuracy for different types of classifiers. We demonstrate that our complete approach can be integrated into a single patient session, including the use of the classifier for functional orthosis-assisted tasks. To the best of our knowledge, this is the first time an intent classifier trained partially on synthetic data has been deployed for functional control of an orthosis by a stroke survivor.
Index Terms:
Generative AI, Synthetic Data, Rehabilitation Robotics, Prosthetics and Exoskeletons, Wearable Robotics.
@@ I Introduction
Wearable rehabilitation robots continuously interact with human patients and must constantly make decisions on when and how to provide motor assistance. Machine learning methods are being applied increasingly to mediate this stream of information. However, many current learning methods, which have revolutionized domains such as vision or language, rely on the availability of large training datasets. Compared to these domains, applications for learning on wearable robots are faced with a tremendous scarcity of both raw data and reliable ground truth labels [1, 2, 3].
One such example is intent inferral, or the process by which a robotic orthosis or prosthesis collects a set of biosignals from the user, and uses them to infer the activity that the user intends to perform, so it can provide physical assistance at the right moment. In the case of a hand orthosis for stroke survivors developed in our lab [4, 5] (Fig. 1), the device can use forearm electromyographic (EMG) data to predict when the user is trying to open the hand, and provide assistance to overcome the muscle spasticity. Widely considered to be a key problem in assistive and rehabilitative robotics [6], an effective intent inferral mechanism can be an intuitive way to control a robotic device.
@@
Like many applications in the assistive and rehabilitation devices domain, a fundamental challenge in intent inferral for stroke is the difficulty of collecting training data. The variation in EMG signals across conditions, sessions, and subjects makes the challenge even more pronounced. Firstly, EMG signal presents different patterns across subjects for the same intent due to variations in neuromuscular control impairments [7, 8]. In addition, defining a use session as a single use of the device between donning and doffing, even for the same subject, the muscle tone and spasticity can vary across different sessions [3]. Furthermore, the signals are non-stationary and could change over time within a single session, depending on the use conditions, such as the hand position and whether the motor is engaged and providing active grasping assistance [1, 2]. Due to such variation, intent classifiers trained for a specific condition/session/subject, do not generalize well, and classical solutions often require tedious data collection on every new condition/session/subject, introducing a significant burden on the participants.
In this work, we aim to reduce the burden of data collection from stroke subjects by generating synthetic data. We propose ChatEMG, an autoregressive generative model that understands the broad behavior of forearm EMG signals from a corpus of offline data across different stroke subjects and then can generate personalized (i.e., condition-, session-, and subject-specific) synthetic samples conditioned on prompts (i.e., a given sequence of EMG signals) sampled from a very limited dataset of a new condition, session, or subject.
@@ II Related Work
There are many previous works that attempt to infer the activity intents of disabled-bodied subjects using EMG signals with machine learning. These predicted intents are often sent to control wearable assistive and rehabilitative devices. A majority of works [1, 9, 10, 11] use supervised learning to control prostheses or orthoses, where the model is trained only on an initial dataset and then used during a longer session. However, it is not easy to collect labeled data on stroke subjects, and a new dataset needs to be collected for every new session after donning and doffing the device. Recent works [2, 12, 13, 14] have been exploring the semi-supervised learning paradigm to control wearable devices with EMG signals. This paradigm leverages unlabeled data to make the intent inferral algorithm more robust to change from the input signals within a single session. It still relies on some heuristics to “label” the unlabeled data, which can be inaccurate from time to time.
@@
The lack of accurate and reliable data is not specific to intent inferral, but to the general machine learning research in the medical community. Similar to ours, there is some previous work that studies generative models for synthetic data generation. A majority of them [15, 16, 17, 18, 19, 20, 21] use Generative Adversarial Network (GAN) [22] and its variants.
GAN and its variants are not autoregressive and can only generate a fixed length of signals from random seeds. In comparison, ChatEMG can generate an unlimited sequence of EMG signals conditioned on EMG prompts, which are also of arbitrary length. This is extremely useful in applications where personalization is necessary. Stroke subjects exhibit different hand functionality and ChatEMG allows us to bias data generation via prompting.
Most similar to ours, Bird et al. [23] train a GPT-2 model to generate EMG signals for hand open/close classification for healthy subjects. However, their method also does not allow conditioning on prompts, and they only study random forest classifiers with healthy subjects. We specifically focus here on stroke patients whose intents are much more challenging to infer and the signal variations across subjects make conditional data generation from prompts critical.
@@ III Overview
The most direct way to achieve this goal is to collect a set of training data (EMG signals) labeled with ground truth intent. This can be done by instructing the patient to attempt one of the three hand movements of interest, while simultaneously recording EMG data and labeling it with the prescribed intent. Once enough labeled data is collected for each of the three possible intent classes, we can train a classifier to distinguish between them.
However, this traditional approach suffers from two key limitations: (1) The process of collecting labeled training data is burdensome and time-consuming for both the patient and the experimenter. It uses up precious session time and also fatigues the patients, leading to increased muscle tone and spasticity. (2) EMG signals exhibit significant variations between different conditions, sessions, and subjects. Training data collected in one condition/session/subject is unlikely to apply to a different one, leading to very poor generalization performance of the classifier.
@@ IV ChatEMG Generative Models
The role of a ChatEMG model is to take in a sequence of EMG signals as input and predict the next EMG signal as output, where one signal consists of the 8-channel data from our EMG armband. We use the Myo armband from the Thalmic Labs, which has 8 electrodes covering the forearm and collects signals at 100Hz. Such a model can then be used autoregressively to generate synthetic data of arbitrary length, all conditioned on the given prompt. We note that this approach is similar in concept to language models such as ChatGPT [24], capable of generating text in response to a text prompt. However, ChatEMG operates on the “language” of EMG data, hence its chosen moniker.
We train one ChatEMG model on data corresponding to each user intent. The goal is then for each of these models to generate synthetic data corresponding to the intent on which it has been trained. While each of these models is trained on different data, their architecture is the same.
@@ IV-A Architecture
One key consideration for designing the ChatEMG architecture is its ability to generate diverse output given the same input signals. As a result, the output of the model is sampled from a probability distribution over a discrete vocabulary, and during training, we compute the cross-entropy loss between the predicted distribution and the one-hot label. Furthermore, we can generate an arbitrary number of “likely” next signals simply by repeating the sampling process. Each sequence can then be continued in an autoregressive fashion. This means that one generative model can use a single prompt to generate an arbitrary number of likely next sequences, each of arbitrary length. This is an important feature of our model, given that we will need large amounts of synthetic data in order to train downstream classifiers [^1] As an alternative, we also tried a variational regression setup that predicts a mean and variance for the next signal, and used the reparameterization trick to backpropagate the loss, which also allows sampling an arbitrary number of completions for a single prompt. However, we found empirically that the balance between the KL (Kullback-Leibler) loss and the reconstruction loss is hard to tune, and its generated signals are not as good as the classification version..
However, this architecture requires a discrete vocabulary of finite size, which is challenging given that our raw data consists of an 8-channel EMG signal. For preprocessing, we smooth out the data using a median filter of size 9 and then bin and clip each channel of the EMG signal to be an integer between 0 and 1000. Under this range, if we model the whole 8-channel signal as one “token” in our vocabulary, then our vocabulary size becomes $1000^{8}$, which is too large to predict a probability distribution and sample from. As a result, ChatEMG always predicts the next EMG value for only the first channel, making the vocabulary size 1000.
@@ IV-B Modelling Inter-channel Relationship
Although our model predicts one channel at a time, it still considers the interchannel relationship. Our EMG armband has 8 electrodes surrounding the forearm, and each electrode covers a particular area of the forearm muscle. The interchannel relationship across different electrodes can be useful information to leverage for understanding user intent.
@@
Our model is trained with EMG sequences of length $t=256$ (2.56 seconds at 100Hz). The dimension of the model output is $256\times 1$, and they are the 256 predicted EMG values for the next time step of the first channel. During training, we also augment the input signals by rotating the channels seven times to simulate the rotation of the electrodes. This data augmentation strategy enables ChatEMG to be invariant with channel rotation. During data generation, we sample EMG prompts of length 150 (corresponding to 1.5s) from the very limited dataset of the new condition, session, or subject and use ChatEMG to autoregressively complete the rest of the signal to a length of 256, which is the time-series length that our classification algorithms take.
@@ V Experimental Setup
The ultimate goal of orthosis is to provide meaningful functional assistance for stroke subjects, enabled by intent inferral. Thus, we examine whether the generated synthetic samples by ChatEMG can improve intent inferral accuracy.
@@ V-A Subjects
We performed experiments with 5 chronic stroke survivors having hemiparesis and moderate muscle tone: Modified Ashworth Scale (MAS) scores $\leq$ 2 in the upper extremity. Our MAS criteria exclude subjects whose fingers are difficult to move passively — fingers with more severe spasticity cannot be quickly extended with external forces without increasing muscle tone and risking damage to the joints. Our participants can fully close their hands but are unable to completely extend their fingers without assistance. The passive range of motion in the fingers is within functional limits. Testing was approved by the Columbia University Institutional Review Board (IRB-AAAS8104) and was performed in a clinical setting under the supervision of an occupational therapist.
@@ V-B Data Collection Protocol
For each stroke subject, we collect data from two sessions on two different days using the following protocol. A session is defined to be a single use of the device between donning and doffing the device. We intentionally keep the two sessions for each subject at least one week apart to better study the variation of the EMG signals across different days. For each session, we collect data under four different conditions: 1) with the arm resting on a table and the orthosis motor off {arm on table, motor off}, 2) with the arm resting on a table and the orthosis motor on, providing active grasp assistance {arm on table, motor on}, 3) with the arm raised above the table and the orthosis motor off {arm off table, motor off}, and 4) with the arm raised above the table and the orthosis motor on {arm off table, motor on}.
@@ V-C Assessment Scenarios
We create different assessment scenarios (listed below) that simulate different use cases of ChatEMG, by selecting different training recordings (recordings used to train ChatEMG) and intent inferral recordings (recordings used to perform intent inferral evaluation). These scenarios evaluate how well ChatEMG generalizes and adapts to new conditions/sessions/subjects not seen in its training recordings.
@@ V-C3 Subject Adaptation
This scenario simulates onboarding new subjects. We conduct five separate experiments, each one simulating the onboarding of one holdout subject, given that we have seen the other four. In each experiment, we train ChatEMG using all the recordings from the other subjects (including both sessions), and the intent inferral recordings are those of the holdout subject. When adapting to a new subject, it is also implicitly adapting to a new session. However, in our session adaption experiments, we assume it is a different session of a previously seen subject.
We use a subset of the training recordings (around 314K samples of size 256 by 8) to train ChatEMG, and we use the remaining training recordings (around 204K samples) as the validation set. We early stop the generative training before the validation loss increases to avoid overfitting.
@@ V-D Intent Inferral Classifiers
ChatEMG is classifier-agnostic, and the generated synthetic samples can be integrated with the training set of any classifiers. We study three types of classifiers: linear discriminant analysis (LDA), random forests (RF), and Transformer. They are popular in the biomedical literature and cover both classic machine-learning algorithms and high-capacity neural networks. We feed into each classifier a time series of length 256 (2.56s), in the shape of 256 by 8. The EMG signals are flattened into a single vector for LDA and RF. We use a single 4-head attention block followed by a 3-layer multilayer perception (MLP) for the Transformer classifier. We perform the same preprocessing techniques as training the ChatEMG model, and as an additional step, we normalize the EMG signals into the range of $[-1,1]$.
@@ V-E Baselines
The intent inferral evaluation is done on individual intent inferral recordings. For each recording, we assume only a small support set (i.e., the first open-relax-close motion of the recording) is available for training the classifier. The support set simulates the limited new training samples from a new condition, session, or subject, denoted by $\mathcal{D}_{\text{new}}^{\text{synth}}$ in Sec. III. We then test the classifier’s accuracy using the query set (i.e., the second and third open-relax-close motions).